OpenStudy (anonymous):

(f-g)(x) (f-x)(g-x) =2x-7-x-2 =?

OpenStudy (mimi_x3):

= x - 9

OpenStudy (anonymous):

=(2x-x) + (-7 - 2) =x - 9